International Journal of Modern Physics C Status Report — January 16, 1998

(P2) “Heavy Quark Potential in Lattice QCD — A Review of Recent Progresses”

Authors: H.-Q. Ding Current Status: Vol. 2, No. 2, pp. 637–658

(P3) “An Efficient Method to Calculate Field Theories with Dynamical Fermions”

Authors: Marcia G. Do Amaral, M. Kischinhevsky, C. Aragao de Carvalho and F. L. Teixeira Current Status: Vol. 2, No. 2, pp. 561–600

(P4) “The 2D Heisenberg Antiferromagnet in High-Tc Superconductivity, A Review of Numerical Techniques

and Results” Authors: T. Barnes Current Status: Vol. 2, No. 2, pp. 659–709

(P5) “Incorporating Knowledge into Simulated Annealing”

Authors: Xin Yao Current Status: Referee comments to authors (4/17/91).

(P6) “Computational Ecosystems in a Changing Environment”

Authors: Natalie Glance, Tad Hogg and Bernardo A. Huberman Current Status: Vol. 2, No. 3, pp. 735–744

(P7) “Physics Goals of the QCD Teraflop Project”

Authors: S. Aoki, et al. (submitted by Carleton DeTar) Current Status: Vol. 2, No. 4, pp. 829–948

(P8) “Calculating Fractal Dimensions”

Authors: S. Goshen and R. Thieberger Current Status: Vol. 3, No. 2, pp. 267–278 (P10) “On Quantum Complexity and the Fundamental Limits of Computational Complexity”

Authors: J. D. Brasher, C. F. Hester, and H. J. Caulfield Current Status: Referee comments to authors (12/2/91).

(P11) “Mathematical Formulations and Computer Enumerations of Polymer Models with Loops” Authors: Hwa A. Lim Current Status: Vol. 3, No. 2, pp. 385–446

(P12) “GISMO: An Object-Oriented Program for High-Energy Physics Event Simulation and Reconstruction” Authors: W. B. Atwood, T. Burnett, R. Cailliau, D. R. Myers, and K. M. Storr Current Status: Vol. 3, No. 3, pp. 459–478

(P13) “Finite-Size Scaling at Critical Points with Spatial Anisotropies”

Authors: Kwan-Tai Leung Current Status: Vol. 3, No. 2, pp. 367–384

(P14) “Monte Carlo Simulation of the Scattering Contributions to the XRF Line”

Authors: Hector Jorge Sanchez and Renzo Sartori Current Status: Rejected (11/2/92).

(P15) “Parallel Acceleration Algorithm for Spin Models”

Authors: G. V. Bhanot and S. Adler Current Status: Vol. 3, No. 4, pp. 605–610

(P16) “Hierarchical Tree-Structures as Adaptive Meshes”

Authors: David J. Edelsohn Current Status: Vol. 4, No. 5,pp. 909–917

(P17) “Modelling and Simulation of Ergodic Data”

Authors: N. Faghih Current Status: Rejected (11/2/92).

(P18) “Estimating Eigenvalue Spectra of Large Matrices”

Authors: R. N. Silver, H. Roder, J. A. Brueggeman Current Status: Paper to Publisher (4/12/94).

(P19) “Spectrum of Text Information Content in the RNA Sequence of the Vesicular Stomatitis Virus” Authors: Alexander A. Filyukov Current Status:

(P20) “A Graphics Interface for Interpreting and Analyzing Inherently Two-Dimensional Images” Authors: Robert J. McKee Current Status: Rejected (7/31/92).

(P21) “Multicanonical Spin Glass Simulations”

Authors: Bernd A. Berg and Tarik Celik Current Status: Vol. 3, No. 6, pp. 1251–1274

(P22) “Glueball and Meson Dispersion Relations in Compact U(1)_{2+1} Lattice Gauge Theory” Authors: A. M. Chaara, H. Kroger, K. J. M. Moriarty, and J. Potvin Current Status: Vol. 4, No. 5, pp. 919–945

(P23) “Rainbows in K_{i η} (ξ)”

Authors: M. S. Hussein and M. P. Pato Current Status: Rejected (2/3/93).

(P24) “V Cycle Dynamical Exponent of the Multi-Scale Update Algorithm for the 2-d XY Model” Authors: S. L. Adler and G. V. Bhanot Current Status: Vol. 4, No. 5, pp. 947–954

(P25) “Spline Techniques for Solving Relativistic Conservation Equations”

Authors: D. J. Dean, C. Bottcher, M. R. Strayer, and G. Gatoff Current Status: Vol. 4. No. 4, pp. 723–747

(P26) “Investigation of the 3-D State, 2-D Potts Model with One Defect Line of Couplings by Density of States

Monte Carlo Method” Authors: Géza Ódor Current Status: Vol. 3, No. 6, pp. 1195 (P27) “A Numerical Implementation of the Dirac Equation on a Hypercube Multicomputer”

Authors: J. C. Wells, A. S. Umar, V. E. Oberacker, C. Bottcher, M. R. Strayer, J.-S. Wu, J. Drake and P. Flanery Current Status: Vol. 4, No. 3, pp. 459–592

(P28) “On the Determinant and Eigenvalues of the Autocorrelation Matrices”

Authors: Nezameddin Faghih Current Status: Rejected (2/3/93).

(P29) “New SIMD Algorithms for Cluster Labeling on Parallel Computers”

Authors: John Apostolakis, Paul Coddington, and Enzo Marinari Current Status: Vol. 4, No. 4, pp. 749–763

(P30) “A Hybrid Monte Carlo Study of the Su-Schrieffer-Heeger Model”

Authors: Marcia G. Do Amaral and C. Aragao de Carvalho Current Status: Paper submitted (11/3/92). Paper to Publisher (9/25/93).

(P31) “Development of a Gaussian Hypergeometric Function Code in Complex Domains” Authors: Yupai P. Hsu Current Status: Vol. 4, No. 4, pp. 805–840

(P32) “Projector Approximation and Quantum Monte Carlo”

Authors: R. M. Fye Current Status: Revised paper received (10/12/93). Paper to Publisher (12/10/93).

(P33) “Automatic Calculations in High Energy Physics by Grace/Chanel and CompHEP” Authors: E. E. Boos, M. N. Dubinin, V. A. Ilyin, A. E. Pukhov, S. A. Shichanin, T. Kaneko, S. Kawabata, Y. Kurihara, Y. Shimizu, and H. Tanaka Current Status: Revised paper received (12/10/93). Paper to Publisher (12/10/93).

(P34) “Autoregressive Genomic Walks”

Authors: B. G. Nguyen, D. J. Cork, D. Wu, T. Grace, and J. Roberge Current Status: Referee comments to authors (7/2/93).

(P35) “Compressibility Effects in Modeling Turbulent High Speed Mixing Layers”

Authors: Cheryl L. Sellers and Suresh Chandra Current Status: Referee comments to authors (8/10/94).

(P36) “Multigrid Methods for the Computation of Propagators in Gauge Fields”

Authors: Thomas Kalkreuter Current Status: Paper to Publisher (3/12/94).

(P37) “Parallel Implementation of 3 + 1-Dimensional Relativistic Hydrodynamics” Authors: D. J. Dean, C. Bottcher, M. R. Strayer, and J. C. Wells Current Status: Vol. 4, No. 5, pp. 1023–1040

(P38) “Toward a Theory of the Physics of the Brain”

Authors: Richard T. Gordon Current Status: Rejected (1/17/94).

(P39) “Spectral Properties of Derivative Operators in the Basis-Spline Collocation Method” Authors: J. C. Wells, V. E. Oberracker, M. R. Strayer, and A. S. Umar Current Status: Paper submitted (8/22/93). Paper to Publisher (10/20/94).

(P40) “Analysis of Random Number Generators using Monte Carlo Simulation”

Authors: P. D. Coddington Current Status: Paper submitted (9/13/93). Paper to Publisher (9/29/93).

(P41) “Absorbing Computer Aided Design Geometries into the GEANT Simulation Program” Authors: S. Youssef, P. Dragovitsch, and J. Womersley Current Status: Paper submitted (9/27/93). Paper to Publisher (10/28/93).

(P42) “Topology Dependence in Lattice Simulations of Non-linear PDEs on a MIMD Computer” Authors: P. Valin, B. Goulard, and M. Sanielevici Current Status: Revised paper received (7/15/94). Paper to Publisher (8/10/94).

(P43) “Full QCD on APE100 Machines”

Authors: S. Antonelli, M. Bellacci, A. Donini, and R. Sarno Current Status: Revised paper received (7/15/94). Paper to Publisher (8/10/94).

(P44) “Vector Hysteresis Modeling”

Authors: Alexandru Stancu and Constantin Papusoi Current Status: Referee comments to authors (1/31/95).

(P45) “Parallel Wolff Cluster Algorithms”

Authors: S. Bae and S. H. Ko Current Status: Paper to Publisher (3/5/95).

(P46) “The Mathematical Structures of Hierarchical Algorithms”

Authors: Zheng-Yao Su and Geoffrey C. Fox Current Status: Referee comments to authors (8/5/94).

(P47) “Monte Carlo Simulation of the DC Size Effect in Thin Films”

Authors: R. I. Boughton and Zhiping Feng Current Status: Paper to Publisher (3/5/95).

(P48) “Massively Parallel Supercomputing for Electromagnetic Modeling” Authors: C. Y. Pan, Thomas J. Timmerman, and H. Q. Lin Current Status: Referee comments to authors (3/5/95).

(P49) “Applying Tabu Search to the Two-Dimensional Ising Spin Glass” Authors: Manuel Laguna and Pablo Laguna Current Status: Paper to Publisher (12/12/94).

(P50) “Electrophoretic Classification of Topologically Nontrivial Macromolecules” Authors: Elena Kenderova and Hwa A. Lim Current Status: Referee comments to authors (1/17/96).

(P51) “Frequency Locking and Synchronization in 2-D Modulated Maps” Authors: Sangeeta D. Gadre and V. S. Varma Current Status: Referee comments to authors (2/15/96).

(P52) “Projection Monte Carlo Methods: An Algorithmic Analysis”

Authors: Nicolas J. Cerf and Olivier C. Martin Current Status: Revised Paper Received (8/19/95). Paper to Publisher (9/11/95).

(P53) “Primality Test via Quantum Factorization”

Authors: H. F. Chau and H.-K. Lo Current Status: Paper to Publisher (11/11/96).

(P54) “Phase Diagram of a Two Dimensional Dilute Binary Alloy”

authors: G. K. Khalil, K. Yaldram, and A. Sadiq Current Status: Paper to Publisher (11/11/96).

(P55) “Boundary Integral Method for Stationary States of Two-dimensional Quantum Systems” Authors: Ioan Kosztin and Klaus Schulten Current Status: Paper to Publisher (4/25/97).

(P56) “Molecular Orbitals Calculation on LiH with Algebraic Treatment of the Integrals” Authors: A. F. M. Righi, and C. A. Kuhnen Current Status: Paper to Publisher (7/30/97).

(P57) “Exploiting Problem Structure as a Search Heuristic”

Authors: Tad Hogg Current Status: Paper to Publisher (12/1/97).

(P58) “Parallel Invaded Cluster Algorithm for the Ising Model”

Authors: Y. S. Choi, and J. Machta Current Status: Referee comments to authors (1/16/98).